A Pipe Fills Up A Tank In 8 Mins. Another Pipe Drains 6 Kilo Liters Of Water In 1 Min. If Both The Pipes Of The Tank Are Opened At Same Time, Then The Tank Gets Filled Up In 20 Mins. What Is The Capacity Of Tank In Kilo Liters?

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Let volume of tank be k

A speed = k/8
B speed = 6

Given that ( k/8. -6) *20 = k

3k/2 = 120 kilo liter
k = 80 kilo liter

Regards